# Post-hoc Study of Climate Microtargeting on Social Media Ads with LLMs: Thematic Insights and Fairness Evaluation

### v1

@@ Abstract.
Climate change communication on social media increasingly employs microtargeting strategies to effectively reach and influence specific demographic groups. This study presents a post-hoc analysis of microtargeting practices within climate campaigns by leveraging large language models (LLMs) to examine Facebook advertisements. Our analysis focuses on two key aspects: demographic targeting and fairness. We evaluate the ability of LLMs to accurately predict the intended demographic targets, such as gender and age group, achieving an overall accuracy of 88.55%88.$5$. Furthermore, we instruct the LLMs to generate explanations for their classifications, providing transparent reasoning behind each decision. These explanations reveal the specific thematic elements used to engage different demographic segments, highlighting distinct strategies tailored to various audiences. Our findings show that young adults are primarily targeted through messages emphasizing activism and environmental consciousness, while women are engaged through themes related to caregiving roles and social advocacy. In addition to evaluating the effectiveness of LLMs in detecting microtargeted messaging, we conduct a comprehensive fairness analysis to identify potential biases in model predictions. We assess disparities in accuracy and error rates across demographic groups using established fairness metrics such as Demographic Parity, Equal Opportunity, and Predictive Equality. Our findings indicate that while LLMs perform well overall, certain biases exist, particularly in the classification of senior citizens and male audiences. The analysis of thematic explanations uncovers recurring patterns in messaging strategies tailored to various demographic groups, while the fairness analysis underscores the need for more inclusive and unbiased targeting methods. By showcasing the efficacy of LLMs in dissecting and explaining targeted communication strategies and by highlighting fairness concerns, this study provides a valuable framework for future research aimed at enhancing transparency, accountability, and inclusivity in social media-driven climate campaigns.
microtargeting, climate campaigns, social media, demographic targeting, fairness, thematic analysis, large language models
journalyear: 2024copyright: rightsretainedconference: ACM Conference; 2024; N, NN, NNNbooktitle: ACM Conference, 2024, N, NN, NNNccs: Information systems Online advertisingccs: Computing methodologies Natural language processingccs: Applied computing Law, social and behavioral sciences
1. Introduction
@@
Climate change represents one of the most pressing global challenges of the 21st century, necessitating widespread public awareness and engagement to drive meaningful environmental action (Moritz and Agudo, 2013; Dessler and Theater, 1995). As traditional media channels evolve, social media has emerged as a pivotal platform for climate communication, enabling organizations, activists, and policymakers to disseminate information, mobilize support, and influence public discourse on environmental issues (Nosek, 2020; Hestres and Hopke, 2017; Adger et al., 2003). The interactive and targeted nature of social media advertising allows for the customization of messages to resonate with specific audiences, thereby enhancing the effectiveness of communication strategies aimed at fostering climate awareness and behavioral change (Bloomfield and Tillery, 2019; Walter et al., 2018; Stoddart et al., 2016).
@@
In recent years, the utilization of microtargeting strategies in social media campaigns has gained significant traction. Microtargeting involves the precise tailoring of messages to distinct demographic segments based on factors such as age, gender, location, and interests (Prummer, 2020; Hersh, 2015; Barbu, 2014). This approach leverages vast amounts of user data to craft personalized content that is more likely to engage and persuade individual users. In the context of climate communication, microtargeting can enhance the relevance and impact of messages, potentially leading to greater public engagement and support for environmental initiatives. However, the sophistication of these strategies also raises critical questions about the transparency, ethical implications, and overall effectiveness of targeted climate messaging (Islam and Goldwasser, 2024a; Islam et al., 2023b).
Figure 1. Example of climate microtargeting.
@@
Figure 1 illustrates the targeted climate advertisements on social media, with a specific focus on demographic targeting. The first ad source targets young adults aged 18-24, with a message encouraging action against climate change and inviting them to become TCI Ambassadors. The second ad source is tailored to a female audience with a message emphasizing the importance of reducing carbon pollution from home electricity and making a pledge for their children’s future. The third ad source targets a male audience, focusing on the benefits of using clean fuel like Ohio Propane for heating rural homes and fueling appliances.
@@
Despite the growing prevalence of microtargeting in climate campaigns, there remains a limited understanding of the specific techniques and linguistic patterns employed to engage different demographic groups. Traditionally, thematic analysis (TA) has been the preferred method for this task, functioning as a qualitative research technique that focuses on identifying patterns, where themes that emerge from the data drive further analysis (Roberts et al., 2019; Vaismoradi et al., 2016, 2013; Braun and Clarke, 2012, 2006; Tuckett, 2005). However, with the surge of textual data in the digital era, there is an increasing trend towards using computational methods for text analysis (Sivarajah et al., 2017; Jagadish et al., 2014; Fan et al., 2014). TA method often falls short in capturing the nuanced and context-dependent nature of targeted communication. This gap highlights the need for advanced analytical tools that can dissect and interpret the complex language and strategies used in microtargeted climate advertisements. Large language models (LLMs) (Brown et al., 2020), with their robust natural language processing (NLP) capabilities, offer a promising solution to this challenge. In this paper, we investigate whether the newly emerged paradigm in NLP; zero-shot prompting of LLMs (Brown et al., 2020) and the practice of providing explanations of answers — is better equipped to address those challenges.
Explanations are fundamental to human learning (Ahn et al., 1992), as they underscore task principles that facilitate broad generalizations (Lombrozo and Carey, 2006; Lombrozo, 2006). Consider the example text ( “The turbines that provide clean wind energy also create new habitats for fish, keeping the ecosystem healthy.” ) from Figure 2(a) for gender identification task. An explanation can elaborate on a brief answer (e.g., male) by connecting it to the broader reasoning process necessary to solve the problem (e.g., “Emphasizing fish habitats may appeal more to men interested in fishing and environmental conservation.”). Thus, explanations enhance understanding by demonstrating how task principles connect questions to their answers.
While LLMs offer powerful capabilities for analyzing and generating text, their widespread use has also highlighted significant challenges related to fairness (Kumar et al., 2023; Li et al., 2023b, a; Sag, 2023) and bias (Lin et al., 2024; Kotek et al., 2023; Fang et al., 2024; Urman and Makhortykh, 2023; Esiobu et al., 2023). Research has shown that LLMs, like other machine learning (ML) models, can inherit and even amplify biases present in the data they are trained on (Blodgett et al., 2020; Binns, 2018). These biases can manifest in various forms, such as differential accuracy across demographic groups (Dacon and Liu, 2021; Lu et al., 2020; Hitti et al., 2019; D´ıaz et al., 2018), harmful stereotypes (Bianchi et al., 2023; Matthews et al., 2022; Alemany et al., 2022), and discriminatory language patterns (Salguero and Espinilla, 2018). The implications are particularly concerning in high-stakes applications such as healthcare, finance, and social media, where biased outcomes can perpetuate inequities and undermine trust in artificial intelligence (AI) systems (Mehrabi et al., 2021; Buolamwini and Gebru, 2018). Addressing these issues requires fairness evaluations using metrics like Demographic Parity, Equal Opportunity, and Predictive Equality to ensure that models perform equitably across all user groups (Žliobaitė, 2017; Hardt et al., 2016). In the context of this study, examining the fairness of LLMs predictions in demographic targeting is crucial for understanding the microtargeted climate messaging.
In this study, we conduct a post-hoc analysis of climate microtargeting practices on social media by leveraging the power of LLMs (OpenAI’s o1-preview model [^1] https://openai.com/index/learning-to-reason-with-llms/). Post-hoc analysis, typically performed after the main experiment or event, allows us to retrospectively evaluate how effective these campaigns are in targeting specific demographics. Building upon data from previous research by Islam et al. (2023b), we investigate the ability of LLMs to accurately detect targeted messaging based on specific demographic variables, including gender and age group. Additionally, LLMs provide explanations for their classification decisions, offering insights into the thematic and linguistic elements used to engage different audiences. Furthermore, we conduct a comprehensive fairness analysis to identify potential biases in model predictions.
@@
Figure 2. Prompt examples for gender prediction (shown as zero-shot). (a) male, (b) female. Inputs are shown in blue, and outputs are shown in red.
Our exploration leads us to the following research questions (RQ), which are crucial for assessing the potential of LLMs to understand microtargeting patterns and provide deeper and more nuanced insights:
@@
  • RQ1: Given a text, can LLMs predict the targeted demographic of the corresponding text accurately and provide the explanation for the reasoning behind the prediction?
  • RQ2: Which words or phrases are most commonly associated with certain demographics?
  • RQ3: What are the recurrent themes and aspects of explanations provided by LLMs?
@@
  • RQ4: How fair are the LLMs predictions in terms of demographic targeting, and what are the disparities in prediction accuracy and error rates across different demographic groups?
@@ 2. Related Work
The intersection of microtargeting, social media, and climate communication has drawn considerable interest in recent years, with studies examining how digital platforms influence public opinion and engagement (Islam and Goldwasser, 2024b, a; Holder et al., 2023; Islam et al., 2023b). The growing capabilities of social media platforms to deliver personalized messages based on user data have sparked a significant body of research into computational advertising (Abrams et al., 2022; Huh and Malthouse, 2020; Yang et al., 2017; Zhang et al., 2017), demographic targeting (Allen, 2022; Rummo et al., 2020; Ribeiro et al., 2019; Kaspar et al., 2019; Jansen et al., 2013; Jansen and Solomon, 2010), and the broader implications of these practices for societal discourse.
@@
Microtargeting on social media platforms such as Facebook has been widely studied in the context of political campaigns (Islam et al., 2023a; Capozzi et al., 2021; Serrano et al., 2020; Silva et al., 2020; Capozzi et al., 2020) and public health messaging (Islam and Goldwasser, 2022; Silva and Benevenuto, 2021; Mejova and Kalimeri, 2020), revealing both the potential benefits and ethical concerns associated with this practice. In the realm of climate communication, microtargeting can be a powerful tool for enhancing message relevance and impact by tailoring content to the specific values and interests of diverse demographic groups (Islam et al., 2023b). However, the effectiveness and ethical implications of such targeted messaging remain under-explored, particularly in terms of how different demographic groups are engaged and whether biases are present in the targeting strategies.
@@
Recent advances in LLMs have demonstrated their capability for in-context learning (ICL), significantly enhancing their ability to perform tasks traditionally handled by humans (Chowdhery et al., 2023; Kojima et al., 2022; Le Scao et al., 2022; Brown et al., 2020; Gilardi et al., 2023; De Paoli, 2023; Dai et al., 2023; Chiang and Lee, 2023; Ziems et al., 2024). This progress suggests a strong potential for effectively applying LLMs to our specific task. In the realms of qualitative research (QR) and NLP, innovative methods are being explored to integrate LLMs into TA. Researchers have proposed various frameworks, including an LLM-in-the-loop approach (Islam and Goldwasser, 2024b, a; Dai et al., 2023), integrating GPT-3 with expert-designed codebooks (Xiao et al., 2023), and developing collaborative interfaces that utilize LLMs for code generation and support in decision-making processes (Gao et al., 2023). Other recent work has shown that LLMs can benefit from examples that decompose the reasoning process (can be seen as an explanation), leading to an answer (Wei et al., 2022). Despite the impressive capabilities of LLMs, there are concerns about fairness, accountability, and transparency in LLMs predictions which have been highlighted in recent literatures (Anthis et al., 2024; Dai et al., 2024; Kumar et al., 2023; Li et al., 2023b; Wu et al., 2023; Bender et al., 2021), emphasizing the need for rigorous evaluation of biases and disparities in model performance across different groups.
In this paper, we leverage zero-shot capabilities of LLMs to identify targeted demographics and provide explanations of demographic targeting regarding climate-related advertisements on Facebook. Besides, we develop a new set of themes and aspects based on those explanations specifically tailored for analyzing messaging. Additionally, we extend previous research by not only focusing on the accuracy of these predictions but also conducting a comprehensive fairness analysis to identify and address potential biases in the model’s performance.
3. Dataset
@@
We investigate the climate campaigns case study for this work. We work on the corpus of 213722137221372 ads released by Islam et al. (2023b). This dataset includes climate-related English ads on Facebook from the US, spanning from January 202120212021 to January 202220222022. Each ad includes the following attributes: ad ID, title, ad description, ad body, funding entity, spend, impressions, and distribution of impressions broken down by gender (male, female, unknown), age (seven groups), and location down to the state level in the USA. Additional details about the dataset can be found in the original publications.
@@
For this work, we consider two demographic indicators, i.e., gender and age group. We consider two gender categories, i.e., male vs female. Regarding age group, we consider four age group categories, i.e., young adults whose age range is 18-24, early working age group (25-44), late working age group (45-64), and senior citizens (65+). We found approximately 227227227 targeted unique ads; among them, 595959 ads targeting only females and 474747 ads targeting only males. However, we find 252525 ads target only young adults, 828282 ads target only the early working age group, 888 ads target only the late working age group, and 666 ads target only senior citizens.
Figure 3. Prompt template for targeted demographic prediction (shown as zero-shot). (a) gender, (b) age group. Inputs are shown in blue, and outputs are shown in red.
4. Task Definition

### v2

@@ Abstract
Climate change communication on social media increasingly employs microtargeting strategies to effectively reach and influence specific demographic groups. This study presents a post-hoc analysis of microtargeting practices within climate campaigns by leveraging large language models (LLMs) to examine Facebook advertisements. Our analysis focuses on two key aspects: demographic targeting and fairness. We evaluate the ability of LLMs to accurately predict the intended demographic targets, such as gender and age group, achieving an overall accuracy of $88.55\%$. Furthermore, we instruct the LLMs to generate explanations for their classifications, providing transparent reasoning behind each decision. These explanations reveal the specific thematic elements used to engage different demographic segments, highlighting distinct strategies tailored to various audiences. Our findings show that young adults are primarily targeted through messages emphasizing activism and environmental consciousness, while women are engaged through themes related to caregiving roles and social advocacy. In addition to evaluating the effectiveness of LLMs in detecting microtargeted messaging, we conduct a comprehensive fairness analysis to identify potential biases in model predictions. We assess disparities in accuracy and error rates across demographic groups using established fairness metrics such as Demographic Parity, Equal Opportunity, and Predictive Equality. Our findings indicate that while LLMs perform well overall, certain biases exist, particularly in the classification of senior citizens and male audiences. The analysis of thematic explanations uncovers recurring patterns in messaging strategies tailored to various demographic groups, while the fairness analysis underscores the need for more inclusive and unbiased targeting methods. By showcasing the efficacy of LLMs in dissecting and explaining targeted communication strategies and by highlighting fairness concerns, this study provides a valuable framework for future research aimed at enhancing transparency, accountability, and inclusivity in social media-driven climate campaigns.
Post-hoc Study of Climate Microtargeting on Social Media Ads with LLMs: Thematic Insights and Fairness Evaluation
Tunazzina Islam Department of Computer Science Purdue University West Lafayette, IN 47907 islam32@purdue.edu                        Dan Goldwasser Department of Computer Science Purdue University West Lafayette, IN 47907 dgoldwas@purdue.edu
1 Introduction
@@
Climate change represents one of the most pressing global challenges of the 21st century, necessitating widespread public awareness and engagement to drive meaningful environmental action Moritz and Agudo (2013); Dessler and Theater (1995). As traditional media channels evolve, social media has emerged as a pivotal platform for climate communication, enabling organizations, activists, and policymakers to disseminate information, mobilize support, and influence public discourse on environmental issues Nosek (2020); Hestres and Hopke (2017); Adger et al. (2003). The interactive and targeted nature of social media advertising allows for the customization of messages to resonate with specific audiences, thereby enhancing the effectiveness of communication strategies aimed at fostering climate awareness and behavioral change Bloomfield and Tillery (2019); Walter et al. (2018); Stoddart et al. (2016).
@@
In recent years, the utilization of microtargeting strategies in social media campaigns has gained significant traction Eldar and Hidir (2025); Islam (2025b); Nistor (2024). Microtargeting involves the precise tailoring of messages to distinct demographic segments based on factors such as age, gender, location, and interests Islam (2025a); Prummer (2020); Hersh (2015); Barbu (2014). This approach leverages vast amounts of user data to craft personalized content that is more likely to engage and persuade individual users. In the context of climate communication, microtargeting can enhance the relevance and impact of messages, potentially leading to greater public engagement and support for environmental initiatives. However, the sophistication of these strategies also raises critical questions about the transparency, ethical implications, and overall effectiveness of targeted climate messaging Islam and Goldwasser (2024a); Islam et al. (2023b).
Figure 1: Example of climate microtargeting.
@@
Figure 1 illustrates the targeted climate advertisements on social media, with a specific focus on demographic targeting. The first ad source targets young adults aged 18-24, with a message encouraging action against climate change and inviting them to become The Climate Initiative (TCI) [^1] https://theclimateinitiative.org/) Ambassadors. The second ad source is tailored to a female audience with a message emphasizing the importance of reducing carbon pollution from home electricity and making a pledge for their children’s future. The third ad source targets a male audience, focusing on the benefits of using clean fuel like Ohio Propane for heating rural homes and fueling appliances.
@@
Despite the growing prevalence of microtargeting in climate campaigns, there remains a limited understanding of the specific techniques and linguistic patterns employed to engage different demographic groups. Traditional method Roberts et al. (2019); Vaismoradi et al. (2016, 2013); Braun and Clarke (2012, 2006); Tuckett (2005) often falls short in capturing the nuanced and context-dependent nature of targeted communication. This gap highlights the need for advanced analytical tools that can dissect and interpret the complex language and strategies used in microtargeted climate advertisements. Large language models (LLMs) Brown et al. (2020), with their robust natural language processing (NLP) capabilities, offer a promising solution to this challenge. In this paper, we investigate whether the newly emerged paradigm in NLP; zero-shot prompting of LLMs Brown et al. (2020) and the practice of providing explanations of answers — is better equipped to address those challenges.
@@
Figure 2: Prompt examples for gender prediction. (a) male, (b) female. Inputs are shown in blue, outputs in red.
Explanations are fundamental to human learning Ahn et al. (1992), as they underscore task principles that facilitate broad generalizations Lombrozo and Carey (2006); Lombrozo (2006). Consider the example text ( “The turbines that provide clean wind energy also create new habitats for fish, keeping the ecosystem healthy." ) from Figure 2(a) for gender identification task. An explanation can elaborate on a brief answer (e.g., male) by connecting it to the broader reasoning process necessary to solve the problem (e.g., “Emphasizing fish habitats may appeal more to men interested in fishing and environmental conservation.”). Thus, explanations enhance understanding by demonstrating how task principles connect questions to their answers.
While LLMs offer powerful capabilities for analyzing and generating text, their widespread use has also highlighted significant challenges related to fairness Kumar et al. (2023); Li et al. (2023b, a); Sag (2023) and bias Lin et al. (2024); Kotek et al. (2023); Fang et al. (2024); Urman and Makhortykh (2023); Esiobu et al. (2023). Research has shown that LLMs, like other machine learning (ML) models, can inherit and even amplify biases present in the data they are trained on Blodgett et al. (2020); Binns (2018). These biases can manifest in various forms, such as differential accuracy across demographic groups Dacon and Liu (2021); Lu et al. (2020); Hitti et al. (2019); Díaz et al. (2018), harmful stereotypes Bianchi et al. (2023); Matthews et al. (2022); Alemany et al. (2022), and discriminatory language patterns Salguero and Espinilla (2018). The implications are particularly concerning in high-stakes applications such as healthcare, finance, and social media, where biased outcomes can perpetuate inequities and undermine trust in artificial intelligence (AI) systems Mehrabi et al. (2021); Buolamwini and Gebru (2018). Addressing these issues requires fairness evaluations using metrics like Demographic Parity, Equal Opportunity, and Predictive Equality to ensure that models perform equitably across all user groups Žliobaitė (2017); Hardt et al. (2016). In the context of this study, examining the fairness of LLMs predictions in demographic targeting is crucial for understanding the microtargeted climate messaging.
In this study, we conduct a post-hoc analysis of climate microtargeting practices on social media by leveraging the power of LLMs (OpenAI’s o1-preview model [^2] https://openai.com/index/learning-to-reason-with-llms/). Post-hoc analysis, typically performed after the main experiment or event, allows us to retrospectively evaluate how effective these campaigns are in targeting specific demographics. Building upon data from previous research by Islam et al. (2023b); Islam and Goldwasser (2024a), we investigate the ability of LLMs to accurately detect targeted messaging based on specific demographic variables, including gender and age group. Additionally, LLMs provide explanations for their classification decisions, offering insights into the thematic and linguistic elements used to engage different audiences. Furthermore, we conduct a comprehensive fairness analysis to identify potential biases in model predictions. Our exploration leads us to the following research questions (RQ), which are crucial for assessing the potential of LLMs to understand microtargeting patterns and provide deeper and more nuanced insights:
@@
  • RQ1: Given a text, can LLMs identify the targeted demographic of the corresponding text accurately and provide the explanation for the reasoning behind the prediction?
  • RQ2: What are the recurring themes and aspects of explanations provided by LLMs?
@@
  • RQ3: How fair are the LLMs predictions in terms of demographic targeting, and what are the disparities in prediction accuracy and error rates across different demographic groups?
The main idea of this paper is to utilize LLMs as a tool to analyze real-world targeting practices, measuring biases and fairness in actual campaigns. We explore how LLMs can be used post-hoc as third-party tools to analyze patterns in targeted communication, especially when internal platform logic is not transparent. While platforms have white-box access, external stakeholders (researchers, auditors, policymakers) do not. Our method offers an explainable approach to reverse-engineer targeting practices and uncover potential bias or messaging disparities.
@@ 2 Related Work
The intersection of microtargeting, social media, and climate communication has drawn considerable interest in recent years, with studies examining how digital platforms influence public opinion and engagement Islam and Goldwasser (2024b, a); Holder et al. (2023); Islam et al. (2023b). The growing capabilities of social media platforms to deliver personalized messages based on user data have sparked a significant body of research into computational advertising Abrams et al. (2022); Huh and Malthouse (2020); Yang et al. (2017); Zhang et al. (2017), demographic targeting Allen (2022); Rummo et al. (2020); Ribeiro et al. (2019); Kaspar et al. (2019); Jansen et al. (2013); Jansen and Solomon (2010), and the broader implications of these practices for societal discourse.
@@
Microtargeting on social media platforms such as Facebook has been widely studied in the context of political campaigns Islam et al. (2023a); Capozzi et al. (2021); Serrano et al. (2020); Silva et al. (2020); Capozzi et al. (2020) and public health messaging Islam and Goldwasser (2022); Silva and Benevenuto (2021); Mejova and Kalimeri (2020), revealing both the potential benefits and ethical concerns associated with this practice. In the realm of climate communication, microtargeting can be a powerful tool for enhancing message relevance and impact by tailoring content to the specific values and interests of diverse demographic groups Islam et al. (2023b). However, the effectiveness and ethical implications of such targeted messaging remain under-explored, particularly in terms of how different demographic groups are engaged and whether biases are present in the targeting strategies.
@@
Recent advances in LLMs have demonstrated their capability for in-context learning (ICL), significantly enhancing their ability to perform tasks traditionally handled by humans Chowdhery et al. (2023); Kojima et al. (2022); Le Scao et al. (2022); Brown et al. (2020); Gilardi et al. (2023); De Paoli (2023); Dai et al. (2023); Chiang and Lee (2023); Ziems et al. (2024). This progress suggests a strong potential for effectively applying LLMs to our specific task. In the realms of qualitative research (QR) and NLP, innovative methods are being explored to integrate LLMs into Thematic Analysis (TA). Researchers have proposed various frameworks, including an LLM-in-the-loop approach Islam and Goldwasser (2024b, a); Dai et al. (2023), integrating GPT-3 with expert-designed codebooks Xiao et al. (2023), and developing collaborative interfaces that utilize LLMs for code generation and support in decision-making processes Gao et al. (2023). Other recent work has shown that LLMs can benefit from examples that decompose the reasoning process (can be seen as an explanation), leading to an answer Wei et al. (2022). Despite the impressive capabilities of LLMs, there are concerns about fairness, accountability, and transparency in LLMs predictions which have been highlighted in recent literatures Anthis et al. (2024); Dai et al. (2024); Kumar et al. (2023); Li et al. (2023b); Wu et al. (2023); Bender et al. (2021), emphasizing the need for rigorous evaluation of biases and disparities in model performance across different groups. In this paper, we leverage LLMs to identify targeted demographics and provide explanations of demographic targeting regarding climate-related advertisements on Facebook. Besides, we develop a new set of themes and aspects based on those explanations specifically tailored for analyzing messaging. Additionally, we extend previous research by not only focusing on the accuracy of these predictions but also conducting a comprehensive fairness analysis to identify and address potential biases in the model’s performance.
3 Dataset
@@
We investigate the climate campaigns case study for this work. We work on the corpus of $21372$ ads Islam and Goldwasser (2024a) originally released by Islam et al. (2023b). This dataset includes climate-related English ads on Facebook from the US, spanning from January $2021$ to January $2022$. Each ad includes the following attributes: ad ID, title, ad description, ad body, funding entity, spend, impressions, and distribution of impressions broken down by gender (male, female, unknown), age (seven groups), and location down to the state level in the USA. Additional details about the dataset can be found in the original publications.
@@
For this work, we consider two demographic indicators, i.e., gender and age group. We consider two gender categories, i.e., male vs female. Regarding age group, we consider four age group categories, i.e., young adults whose age range is 18-24, early working age group (25-44), late working age group (45-64), and senior citizens (65+). However, Facebook Ad Library API doesn’t provide explicit targeting information beyond ad impressions (views). We focus only on ads viewed exclusively by either males or females; ads viewed by both genders were excluded. Therefore, if an ad is viewed solely by males or solely by females, we consider it as a targeted ad. The same logic is applied to age groups. We do not include any ads that overlap across age categories such as young adults, early working-age individuals, late working-age individuals, and senior citizens. By following this methodology, we narrow down the dataset to $227$ ads. Among them, $59$ ads targeting only females and $47$ ads targeting only males. However, we find $25$ ads target only young adults, $82$ ads target only the early working age group, $8$ ads target only the late working age group, and $6$ ads target only senior citizens.
4 Experimental Setup